OpenStudy (anonymous):

Explain how to convert 5.38 x 10^10 from scientific notation to standard form, then give the answer in standard form

OpenStudy (anonymous):

10^10=10000000000 (10*10*10...*10 ) 10 times 5.38=538 * 10^(-2 ). 538*10^(-2) *10^10 =538 * 10^8 =53,800,000,000. (10^8=10^ (-2) * 10^10